To support the hydrology studies in the Sierra Ladrones Study Basin, a network of moisture potential sensors and temperature sensors wereinstalled in the stream-channel sediments and adjacent soils atvarious locations up thru the watersheds in 1992. Two rain gaugeswere also added up through the watershed gradient to complement therain gauge on the weather station (Met43) at the base of the watershedto provide a better measure of moisture inputs to these watersheds.This file contains data for 1992 to 1994.
To monitor temperature and moisture of stream sediments and adjacent soils up the gradient of a series of small watersheds in the Sierra Ladrones.
An array of soil moisture potential gypsum blocks were placed in the channel sediments up thru the set of small watersheds. The number of sensors used and design of the placement of sensors was different at each site. All probes were initially buried at 20 cm but this depth must surely have changed with time and stream flow. Temperature probes were buried at the same time and the same depth as the gypsum blocks. Tipping bucket rain gauges were installed at sites 23 and 25 to complement the gauge at the Red Tank (Met43) Weather Station
Automated dataloggers with attached soil moisture potential and temperature sensors.
Probes were measured every 30 sec and summary data output at the hour. Precipitation was output at 1-min intervals during precipitation events.
Manufacturer: Campbell Scientific Inc.
Cable Tester Multiplexer SDMX50
TDR probe 30 cm
Aug 14, 1992 - Corrected watershed programs. Found problems w/ soil moisture on 23 and 25 on sensors 3 and 4. Also fixed precipitation program on all dataloggers. Changed out 4th temp sensor because it had been reading too high.
July 31, 1992 - Repaired damage at 23.
Jul 28, 1992 - Repaired damage at 22
Jul.21, 1992 - Storm washed out sensors at 22 and 23
Apr. 17, 1992 - Completed installation of dataloggers. Added 22 at Abeja, Venado Added upper temp and soil moisture sensors at 23; Added precip gages at 23 and 25. D.M.
Mar. 17, 1992 - Installed dataloggers at watersheds. Installed 21 (Above Red Tank),23 (Abeja Seep), 24 (Junction Ensenal and Yucca), 25 (Top of Ensenal). D.M.
Oct 22, 1993 - Sta 23 st3 and sm3 displaced -replaced; Sta 21 st1 indicates probe less deep; Sta 24 st2 indicates probe more shallow; GPS'd in all data loggers and probes - also traced center of channel on trip up and trip down.
07 Oct 1993. - Began dbf for Watershed data
Aug. 30, 1993 - FLOOD!!!!!! 1.85-2.00" of precip at about 1800.
May 11, 1993 - Replaced gypsum soil moisture blocks at Abaja seep - sm2, sm3, sm4, blocks were all eaten away. Reinstalled new ones at almost the same spot as the old. 2nd is only 15-17 cm deep 3rd only about 17 cm as well. 4th (side seep) was moved a little further away from the seep so that it was actually at 20 cm. All 3 are set on top of bedrock and can't be expected to last more than a year if seeps keep flowing. No surface standing water but some of the sediments are still slightly damp.
April 7,1993 - Replaced 2 of the soil moisture probes that had been chewed through by a deviant rodent sm1 and sm2.
Aug 5, 1994 D.M.
25 Found all analog grounds and grounds disconnected- Not sure how long this been the case. SM4 had all wires disconnected; Reattached - nothing reading still but very dry.
23 Precipitation leads loose - had been pulled loose - tail of P1 still in datalogger slot.
June 24, 1994
21 Temperature probe on the far side of the channel was pointing upward towards the surface of the soil. Depth was on the order of 10-15 cm. Soil Moisture block exposed, located probe and reburied the temp probe in the same vicinity.
23 Soil moisture probes at locations 2 (among willows) and 4 (side seep) were located and discovered to be deteriorated. These were removed and new probes were placed and buried. All cables at the pole were retaped to discourage rodent activity.
24 The temp probe on the right hand channel (Yucca) was closer to the surface than normal due to erosion of that channel. Depth was about 15 cm. We reset both probes to 20 cm and verified the decreasing soil temp. At the time of departure, both temp probes were reading similar temps.
file created 12/15/1998 - D.M.
Mar 8, 1995 - Removed 24 datalogger to use for TDR - sensors still in place. Removed precip gauge at 23
Feb 2, 1995 - Removed datalogger @ 21 to use with TDR - sensors still in place
Additional Information on the Data Collection Period
See all Sevilleta Publications